# Algebraic Control: Complete Stable Inversion with Necessary and Sufficient Conditions

### v1

@@ III-D Robustness
where $\epsilon>0$ can be taken arbitrarily small which ensures that the Laplace transform is taken in a way that sidesteps any instantaneous feedback algebraic loop. Consider now the closed-loop configuration depicted in Figure 1, as
$U_{fb}(s)$ $=\Xi_{i}^{\ell}(s)\bar{Y}_{\Delta}(s)=\frac{z^{+}(s)}{z_{d}^{+}(s)}\Xi_{i}^{\ell}(s)Y_{\Delta}(s)$ (53)
$=\frac{z^{+}(s)}{z_{d}^{+}(s)}\Xi_{i}^{\ell}(s)\bigr{(}\operatorname*{proj}_{\operatorname*{Im}(P)}[Y_{\Delta}]+\operatorname*{res}[Y_{\Delta}]\bigl{)}$ (54)
where $\langle\operatorname*{res}[Y_{\Delta}],q_{k}\rangle=0$ implying $\Xi_{i}^{\ell}(s)\operatorname*{res}[Y_{\Delta}]=0$. Thus, the feedback only responds to the part of $Y_{\Delta}$ in $\operatorname*{Im}_{\mathbb{R}(s)}(P)$. Consequently,
$PU_{fb}=\frac{z^{+}(s)}{z_{d}^{+}(s)}\operatorname*{proj}_{\operatorname*{Im}(P)}[Y_{\Delta}]$ (55)
Next, we express the integral equation over $Y_{\Delta}$:
$(I+\Delta_{u})P(s)U_{c}(s)-P(s)U_{c}(s)=\Delta_{u}P(s)U_{c}(s)$ (56)
$=Y_{\Delta}(s)=\underbrace{\Delta_{u}P(s)U_{ff}(s)}_{\begin{subarray}{c}\in\mathcal{RH}_{\infty}\text{ and}\\ \text{ invariant under }u_{fb}\end{subarray}}-\Delta_{u}P(s)U_{fb}(s).$ (57)
then using (53) yields
$\Delta_{u}P(s)U_{ff}(s)-\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\ell}(s)Y_{\Delta}(s)=Y_{\Delta}(s)$
$\Delta_{u}P(s)U_{ff}(s)=\Bigl{(}I+\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\ell}(s)\Bigr{)}Y_{\Delta}(s)$
@@
$\biggl{\|}\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\ell}(s))\biggr{\|}_{\infty}=\alpha_{b}<1$
@@
$\sum_{k=0}^{\infty}\biggl{(}\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\ell}(s)\biggr{)}^{k}$ (58)
Thus, $\Bigl{(}I+\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\ell}(s)\Bigr{)}^{-1}$ exists and stable with $\|Y_{\Delta}\|_{\infty}\leq\frac{\|\Delta_{u}P(s)U_{ff}(s)\|_{\infty}}{1-\alpha_{b}}$. Then, the tracking error is
$Y_{d}(s)-Y_{\Pi}(s)$ $=Y_{d}(s)-Y_{\Delta}(s)-\underbrace{P(s)U_{ff}(s)}_{\triangleq Y_{d}(s)}+P(s)U_{fb}(s)$
$=-Y_{\Delta}+P(s)U_{fb}(s)$
@@
This shows that the tracking error converges to the inverse Laplace transform of the non-cancellable yet stable residual term, thus completing the proof.
@@
Assume $P(s)$ is square, full rank implying $\operatorname*{res}[Y_{\Pi}]=0$, and $Z_{R}\neq\emptyset$, where the control inputs be defined as $U_{ff}(s)=S_{i}(s)K(s)Y_{d}(s)$ and $U_{fb}(s)=S_{i}(s)K(s)\bar{Y}_{\Delta}(s)$. Then (51) becomes $\bigl{(}y_{d}(t)-{y}_{\Pi}(t)\bigr{)}\to 0$.
@@
$\implies\bigl{(}y_{d}(t)-y_{\Pi}(t)\bigr{)}=\mathscr{L}^{-1}\{\text{res}[Y_{\Pi}]\}\ \forall t>0$
Here, the term $\mathscr{L}^{-1}\{\text{res}[.]\}$ corresponds directly to $\varphi$ as described in (9), representing the contribution of inevitable errors. On the other hand, from an implementation perspective, the compactness of the algebraic structures allows for explicitly stating the existence and boundedness conditions and facilitates straightforward solutions of stable inversion. However, when utilizing $\mathcal{H}_{\infty}$-based approximate solutions, we encounter complex, high but finite-order algebraic structures. These structures should be reduced to an acceptable order through balanced model reduction techniques.
@@ IV Numerical Examples
In this section, we present several numerical examples to illustrate the practical application and effectiveness of our proposed approach for model inversion. These examples cover various scenarios demonstrating the broad applicability and flexibility of the proposed method. Lastly, these examples aim to provide an intuitive understanding of the theoretical concepts previously discussed.
@@
Consider the following state-space representation.
$\dot{x}(t)$ $=\begin{bmatrix}-3.565&-0.1919&0.09499\\ 4.418&-1.44&-0.0847\\ -0.3121&0.157&-0.005135\end{bmatrix}x(t)$ (59)
$\quad+\begin{bmatrix}-1.555&0.4156\\ 1.173&0.4576\\ -0.05738&0.1017\end{bmatrix}u(t)$
$y(t)$ $=\begin{bmatrix}-0.2573&0.4612&-1.024\\ 0.2573&-0.4612&1.024\end{bmatrix}x(t)$
where the two-input-two-output system is minimal, and has an invariant zero at $s=1$ (so it is a non-minimum phase). Now define $P(s)\triangleq C(sI-A)^{-1}B$. Then, $\operatorname*{rank}_{\mathbb{R}(s)}(P(s))=1<n_{u}=n_{y}$, leading to a so-called singular MIMO that fits Theorem 1. Next, consider the desired output $Y_{d}(s)\triangleq\begin{bmatrix}{1}/{s}&-{1}/{s}\end{bmatrix}^{T}$. In accordance with Assumption 2, we modify $Y_{d}(s)$ to be $Y_{d}(s)=\frac{s-1}{0.5s+1}\begin{bmatrix}\frac{1}{s}&-\frac{1}{s}\end{bmatrix}^{T}$. This makes $Y_{d}(s)$ exactly trackable such that $\operatorname*{rank}_{\mathbb{R}(s)}\left[P(s)\right]=\operatorname*{rank}_{\mathbb{R}(s)}\left[P(s):Y_{d}(s)\right]=1$. Then the resulting control input $u(t)$ in (59) is
$U(s)=\begin{bmatrix}\frac{2s+6}{s}-\frac{100}{100s+1}\\ 1\end{bmatrix}$ (60)
then to eliminate the Dirac-delta functions resulting from inverse Laplace transformation, multiplying (60) with a suitable low-pass filter and taking the inverse Laplace transformation yields
$u(t)=\mathscr{L}^{-1}\Bigl{\{}U(s)\frac{1}{0.01s+1}\Bigr{\}}=\begin{bmatrix}195e^{-100t}-e^{-0.01t}+6\\ 100e^{-100t}\end{bmatrix}$
The tracking result is given by Fig. 2. It is shown that both outputs track the reference signals which validates the theory where $y_{d}(t)=y(t)$.
@@
[width=0.9]ex1_outs
Figure 2: The output tracking of Example 1 as Theorem 1.
@@
As a comparison to model matching problem in [11], consider the following $2\times 2$, full rank and minimal system that has an invariant zero at $s=1$.
In this case, $\operatorname*{rank}_{\mathbb{R}(s)}(P(s))=2=n_{u}=n_{y}$, which implies that all possible output combinations can be tracked. Given Theorem 3, the modifications introduced by Assumption 2.(2) are unnecessary. Therefore, let’s consider $Y_{d}(s)=\begin{bmatrix}{1}/{s}&{1}/{s}\end{bmatrix}^{T}$, which results in $\operatorname*{rank}_{\mathbb{R}(s)}\left[P(s)\right]=\operatorname*{rank}_{\mathbb{R}(s)}\left[P(s):Y_{d}(s)\right]=2$. We can solve equation (III-C) for (61) with $\gamma=3.6$. As $S_{i}(s)K(s)$ is strictly proper, this is all in line with Theorem 3. From the solution, it can be observed that both Theorem 2 and Lemma 1.(1) are satisfied. Consequently, the resulting $u(t)\triangleq\mathscr{L}^{-1}\{U(s)\}$ in equation (31) is depicted in Fig. 3 and resulting $S_{i}(s)K(s)$ and $T_{i}(s)$ are given by Fig. 4. Compared with the results obtained using the Inner-Outer Factorization method (refer to [11, Fig.8 - black lines]), the proposed method considerably enhances tracking performance (almost at par with the feedback case) in the feedforward scenario.
@@
[width=0.9]ex2_outsandinps.pdf
Figure 3: The output tracking of Example 2 and corresponding inputs as Theorem 3.
[…]
[width=0.9]ex2_SKT.pdf
Figure 4: The singular value plots of $S_{i}(s)K(s)$ and $T_{i}(s)$.
Example 3.
Now, consider the following $n_{y}=2,n_{u}=1$ system which is an example of an underactuated system.
$P(s)=\begin{bmatrix}\frac{s^{2}-5s-50}{s^{2}+3s+2}\\ \frac{s-10}{s^{2}+3s+2}\end{bmatrix}$ (62)
where the system has an invariant zero at $s=10$, thus it is also a non-minimum phase system. Now since choosing $Y_{d}\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$, is much straightforward, we set $Y_{d}$ as
$Y_{d}(s)=\begin{bmatrix}\frac{2}{s^{2}+1}+\frac{8}{s}\\ \frac{1}{s}\end{bmatrix}\notin\operatorname*{Im}_{\mathbb{R}(s)}(P)\text{ s.t.}\operatorname*{res}(Y_{d})\neq 0$ (63)
Then, let’s choose the scalar $P_{p}(s)=\frac{s^{2}-5s-50}{s^{2}+3s+2}$ which also has an invariant zero at $s=10$ and thus no need to augmentation of $z^{+}$, and then solving (III-C) yields a stable $T_{p}(s)$ with $\gamma=2.32$. Also, based on (62), we have
Then using Theorem 4, we have a strictly proper controller resulting in $T_{p}(s)P^{\dagger}(s)$. The reference tracking performance of the form in Theorem 4 shown by Fig. 5 and the design of $T_{p}(s)$ and the boundedness of $u(t)$ are given by Fig. 6. As it can be seen from Fig. 5, the errors are fully in harmony with Definition 1.
[…]
[width=0.9]ex3_outs.pdf
Figure 5: The output tracking of Example 3 as Theorem 4.
[…]
[width=0.9]ex3_inpandT.pdf
Figure 6: The control input and singular values of $T(s)$ of Example 3.
Example 4.
We again consider the same nominal system as in (61) whose invariant zeros are located at $s_{1,2,3}=-10,-0.86,1$ and poles at $s_{1,2,3,4}=-1,-1,-0.5,-3$. Then keep $S_{i}(s)K(s)$ of Example 2 as a controller for $\gamma=3.6$ in Fig. 4. Then, $S_{i}(s)K(s)$ is used both for feedforward and feedback actions in Fig. 1. Then, we start perturbating the system resulting as
${P}_{\Pi}(s)=\begin{bmatrix}\frac{(0.2-s)}{s^{2}+20s+100}&\frac{s^{2}+1.8s-0.4}{s^{2}+8s+16}\\ \frac{0.3}{s+0.1}&\frac{2}{s+2.1}\end{bmatrix}$ (65)
whose invariant zeros are located at $s_{1,2,3,4,5}=-20.4,-6.4,-2.7,-1.2,0.2$ and poles at $s_{1,2,3,4,5,6}=-10,-10,-4,-4,-0.1,-2.1$.
Example 5.
Consider the following state-space representation.
$\dot{x}(t)$ $=\begin{bmatrix}-0.8808&1.876&-1.011&0.8287\\ 0.1876&-0.5055&-0.0869&-0.8554\\ 0&0&-0.9163&0\\ 0&0&0&-0.5108\end{bmatrix}x(t)$ (66)
$\quad+\begin{bmatrix}2.196&-1.785\\ -0.1832&1.813\\ 1.527&0\\ 0&1.277\end{bmatrix}u(t)=Ax(t)+Bu(t)$
$y(t)$ $=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\end{bmatrix}x(t)=Cx(t)$
Equation (66) discrete form with a sample time of $1$ seconds is given by [15], geometric approaches in multivariable control, which is a benchmark in the literature. $C(sI-A)^{-1}B$ is a full-rank, minimal, and square system implying all possible desired output can be achieved whether bounded or unbounded control inputs. Moreover, $Z_{P}=\{-0.2149,0.0925\}$, $Z_{R}=\{0.0925\}$ indicating the system has a dominant RHP zero making the system non-minimum phase. The reference signal is
$y_{d}(t)=\begin{bmatrix}y_{d1}(t)\\ y_{d2}(t)\end{bmatrix}=\begin{cases}[\frac{t}{20}\ 0]^{T}&\text{ for }0\leq t<20\\ [1\ 0]^{T}&\text{ for }t\geq 20\end{cases}$ (67)
$Y_{d}(s)=\begin{bmatrix}(1-e^{-20s})/(20s^{2})\\ 0\end{bmatrix}\text{ s.t. }Y_{d}(0.0925)\neq\begin{bmatrix}0\\ 0\end{bmatrix}$ (68)
Given conditions show that Theorem 3 applies for the feedforward part in Fig.1 with setting $u_{fb}(t)=0$. We can solve equation (III-C) for (66) with a finite $\gamma$ where resulting $S_{i}(s)K(s)$ and $T_{i}(s)$ functions, given by Fig. 7, are strictly proper. This is all in line with Theorem 3.
[…]
[width=0.9]ex6_SKT.pdf
Figure 7: The singular value plots of $S_{i}(s)K(s)$ and $T_{i}(s)$ for Example 5.
After that, the controller $U^{a}(s)$ is discretized by Bilinear transform with $1$ seconds of sample time in order to align the results to the ones in [15, Section 5]. The discrete desired outputs, $y_{d}(t)$, the outputs, $y_{1}(t),y_{2}(t)$-proposed, with respect to the implemented controller $U^{a}(s)$ in this proposed approach, and the outputs, $y_{1}(t),y_{2}(t)$-geometric, under geometric control are given in Fig. 8.
[…]
[width=]ex6_outs.pdf
Figure 8: The output tracking comparison of approximated control in Theorem 3 to geometric approaches in [15].
Finally, the corresponding inputs are given in Fig. 9.
[…]
[width=0.9]ex6_ins.pdf .
Figure 9: The control inputs produced by $U^{a}(s)$
From the results, it is evident that the 20th-degree $U^{a}(s)$ obtained with Theorem 3 and discretized by using Tustin transformation provides almost the same tracking performance as the geometric approach. The most important detail here is that the computation of $U^{a}(s)$ does not require a 100-second preview used in geometric control which opposes causality. Additionally, the structure in (III-C) offers design flexibilities for practitioners to adjust performance, usage, and size for implementation.

### v2

@@ III-D Robustness
where $\epsilon>0$ is chosen as small as to avoid the algebraic loop issue. Consider now the closed-loop configuration depicted in Figure 1, as
$U_{fb}(s)-U_{fb}^{(0)}(s)$ $=\Xi_{a}^{\Gamma}(s){Y}_{\Delta}(s)=\Xi_{i}^{\Gamma}(s)\frac{z^{+}(s)}{z_{d}^{+}(s)}Y_{\Delta}(s)$ (51)
$=\frac{z^{+}(s)}{z_{d}^{+}(s)}\Xi_{i}^{\Gamma}(s)\bigr{(}\operatorname*{proj}_{\operatorname*{Im}(P)}[Y_{\Delta}]+\operatorname*{res}[Y_{\Delta}]\bigl{)}$
where $\langle\operatorname*{res}[Y_{\Delta}],q_{k}\rangle=0$ implying $\Xi_{i}^{\Gamma}(s)\operatorname*{res}[Y_{\Delta}]=0$. Thus, the feedback only responds to the part of $Y_{\Delta}$ in $\operatorname*{Im}_{\mathbb{R}(s)}(P)$. Consequently,
$PU_{fb}=\frac{z^{+}(s)}{z_{d}^{+}(s)}\operatorname*{proj}_{\operatorname*{Im}(P)}[Y_{\Delta}]+PU^{(0)}_{fb}(s)$ (52)
Next, by noting Remark 6, Theorem 3-4, and $P(s),\Delta_{u}\in\mathcal{RH}_{\infty}$, define
$D(s)\triangleq\bigl{[}\Delta_{u}P(s)\bigl{(}U^{(0)}_{ff}(s)-U^{(0)}_{fb}(s)\bigr{)}+Y_{\Pi}^{(0)}(s)+U_{ff}^{(f)}(s)\bigr{]}$
where $Y_{\Pi}^{(0)}(s)=\mathscr{L}^{-1}\{C_{\Pi}e^{A_{\Pi}t}\,x_{0}\}$ s.t. $U^{(0)}_{ff}(s),U^{(f)}_{ff}(s),$ $U^{(0)}_{fb}(s),Y_{\Pi}^{(0)}(s)\in\mathcal{RH}_{\infty}$ are invariant under feedback, thus can be considered as a convergent disturbance acting on the output. Then, the integral equation over $Y_{\Delta}$:
$(I+\Delta_{u})P(s)U_{c}(s)-P(s)U_{c}(s)+Y_{\Pi}^{(0)}(s)=$
$\Delta_{u}P(s)U_{c}(s)+Y_{\Pi}^{(0)}(s)=Y_{\Delta}(s)=D(s)-\Delta_{u}P(s)U_{fb}^{(f)}(s).$
Using (51) yields
$D(s)-\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\Gamma}(s)Y_{\Delta}(s)=Y_{\Delta}(s)$
$D(s)=\Bigl{(}I+\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\Gamma}(s)\Bigr{)}Y_{\Delta}(s)$
@@
$\biggl{\|}\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\Gamma}(s)\biggr{\|}_{\infty}=\alpha_{b}<1$
@@
$\sum_{k=0}^{\infty}\biggl{(}\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\Gamma}(s)\biggr{)}^{k}$ (53)
Thus, $\Bigl{(}I+\frac{z^{+}(s)}{z_{d}^{+}(s)}\Delta_{u}P(s)\Xi_{i}^{\Gamma}(s)\Bigr{)}^{-1}$ exists and stable with $\|Y_{\Delta}\|_{\infty}\leq{\|D(s)\|_{\infty}}/{\left(1-\alpha_{b}\right)}$. Then, the tracking error is
$Y_{d}(s)-Y_{\Pi}(s)$ $=Y_{d}(s)-Y_{\Delta}(s)-P(s)U_{c}(s).$
With $\lim_{s\to 0}sD(s)=0$, Section III-C provides that $P(s)U_{ff}(s)\to Y_{d}(s)$ and thus
$Y_{d}(s)-Y_{\Pi}(s)$ $\to-Y_{\Delta}+P(s)U_{fb}(s)$
@@
This shows that the tracking error converges to the inverse Laplace transform of the non-cancellable yet stable residual term, thus completing the proof. ∎
@@
Assume $P(s)$ is square, full rank implying $\operatorname*{res}[Y_{\Pi}]=0$, and $Z_{R}\neq\emptyset$. Let $U_{ff}(s)=S_{i}(s)K(s)Y_{d}(s)$ and $U_{fb}(s)=S_{i}(s)K(s)\bar{Y}_{\Delta}(s)$. Then, $\bigl{(}y_{d}(t)-{y}_{\Pi}(t)\bigr{)}\to 0$.
@@
$\implies\bigl{(}y_{d}(t)-y_{\Pi}(t)\bigr{)}=\mathscr{L}^{-1}\{\text{res}[Y_{\Pi}]\}$
Here, the term $\mathscr{L}^{-1}\{\text{res}[.]\}$ corresponds directly to $\varphi$ as described in (9), representing the contribution of inevitable errors. In this paper, time delays are not treated as part of the uncertainty; however, the framework can be extended to handle stochastic and delayed systems by leveraging mean-square exponential stability techniques [33]. On the other hand, for implementation, the compactness of the algebraic structures allows for facilitating straightforward solutions. However, when utilizing $\mathcal{H}_{\infty}$-based approximate solutions, we encounter complex, high but finite-order structures where using the balanced model reduction is a solution.
@@ IV Numerical Examples
In this section, we present some numerical examples to illustrate the effectiveness of the proposed approach.
@@
Consider, as in [9], the following $2\times 2$, full rank, and minimal system with $Z_{P}=\{-10,-0.86,1\}$, $Z_{R}={1}$.
In this case, $\operatorname*{rank}_{\mathbb{R}(s)}(P)=2=n_{u}=n_{y}$, which implies that $\operatorname*{res}[Y]=0$. For the desired loop shape $\text{diag}\left[\frac{5}{s},\frac{5}{s}\right]$, we solve (III-C) with $\gamma=3.6$. The approximate inverse, given by $\Xi_{a}^{\Gamma}(s)=S_{i}(s)K(s)$, is shown in Fig. 2.c, consistent with Theorem 3. The resulting control input, $u_{inv}(t)$, is depicted in Fig. 2.b. Compared to the results obtained using the Inner-Outer Factorization (refer to [9, Fig. 8, black lines]), the proposed method significantly improves tracking performance, even using a purely feedforward control scenario.
@@
[width=0.9height=0.095]ex2_SKT2.pdf
Figure 2: The output tracking of Example 1 - (a) - corresponding inputs as Theorem 3 - (b) - approximate inversion - (c) -
Now, consider that the real system $P_{\Pi}(s)$ differs from $P(s)$:
$P_{\Pi}(s)=\begin{bmatrix}\frac{(0.2-s)}{s^{2}+20s+100}&\frac{0.3}{s+0.1}\\ \frac{s^{2}+1.8s-0.4}{s^{2}+8s+16}&\frac{2}{s+2.1}\end{bmatrix},$ (55)
where $\text{Z}_{P_{\Pi}}=\{-20.4,-6.4,-2.7,-1.2,0.2\}\neq\text{Z}_{P}$ and $\text{Z}_{R_{\Pi}}=\{0.2\}\neq\text{Z}_{R}$. Moreover, $P_{\Pi}(s)$ has poles at $s_{1,2,3,4,5,6}=-10,-10,-4,-4,-0.1,-2.1$, which are different from the poles of the nominal system $P(s)$, $s_{1,2,3,4}=-1,-1,-0.5,-3$. According to Theorem 5, or more specifically, Corollary 5.II, to handle the uncertainty, we employ Fig. 1.(b) with $S_{i}(s)K(s)$ by letting $\epsilon_{i}\to 0$ for both inversion blocks. Two simulation, at $t=125^{th}$ sec, a unit step disturbance applies to the output are run. One is with zero initial states for both inverse and forward systems, and the other one tests the proposed method under random initial states as: $x_{0}=[3.1,4.1,-3.7,4.1,1.3,-4],\bar{x}_{0_{1}}=[-0.2,0.6,1.9,1.9,-0.5,1.9,1.8,0.4,1.4,-0.6,0.3,1.7,1.4,1.8],\bar{x}_{0_{2}}=[0.3,-0.9,0.7,0.9,0.4,0.5,0.5,0.5,-0.2,0.3,-0.7,0.4,-1,-0.4,-0.9].$ The reference tracking performance shown by Fig. 3 with a bounded input validating the theory.
[…]
[width=0.8]ex4_outs
Figure 3: The output tracking of Example 1 under uncertainty with zero and arbitrary initial states.
@@
Now, consider the following vectorial system
$P(s)=\begin{bmatrix}\frac{s^{2}-5s-50}{s^{2}+3s+2}&\frac{s-10}{s^{2}+3s+2}\end{bmatrix}^{T}$ (56)
where it has an invariant zero at $s=10$, thus it is a non-minimum phase and underactuated system. Define $Y_{d}$ as
$Y_{d}(s)=\begin{bmatrix}\frac{2}{s^{2}+1}+\frac{8}{s}&\frac{1}{s}\end{bmatrix}^{T}\notin\operatorname*{Im}_{\mathbb{R}(s)}(P)\text{ s.t. }\operatorname*{res}[Y_{d}]\neq 0.$ (57)
Then, $Y(s)=P(s)\Xi_{a}^{\Gamma}(s)Y_{d}(s)$ such that $\operatorname*{res}[Y]\neq 0\impliedby Y_{d}(s)\notin\operatorname*{Im}_{\mathbb{R}(s)}(P)\cap\mathcal{RH}_{\infty}$. Based on Theorem 4, define $P_{p}(s)=\frac{s^{2}-5s-50}{s^{2}+3s+2}$ which also has an invariant zero at $s=10$. Solving (III-C) yields a stable $T_{p}(s)$ with $\gamma=2.32$. Also, based on (56), we have
yielding $T_{p}(s)P^{\dagger}(s)\in\mathcal{RH}_{\infty}$. The reference tracking performance of the form in Theorem 4, the design of $T_{p}(s)$, and the boundedness of $u(t)$ are shown by Fig. 4. As it can be seen from Fig. 4, the errors are fully in harmony with Definition 1.
@@
[width=0.9height=0.1]ex3_outs.pdf
Figure 4: The output tracking of Example 2 - (a) - corresponding inputs as Theorem 4 - (b) - approximate inversion $T_{p}(s)$ (c)